OpenStudy (anonymous):

Let y=safe load in pounds and x= depth in inches for a certain type of rectangular horizontal beam. A constant of proportionality k exists such that y=kx^2. What kind of variation is this relationship?

OpenStudy (anonymous):

exponentially proportional?

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

A: Direct B: Inverse C: Joint D: Combined E: None

OpenStudy (anonymous):

I will have to go for direct, I think.

OpenStudy (anonymous):

Ok. & the next question ask to determine the constant K for a beam with y=800lbs, and x=4 inches.

OpenStudy (anonymous):

Well, that seems straightforward, just plug the numbers into the formula 800 = k * 4^2 -> k = 800/16 or 50

OpenStudy (anonymous):

The next one says: what is a safe laod if x=6 inches. So it would be 800=k*6^2 right?

OpenStudy (anonymous):

Um, no...y is the safe load.

OpenStudy (anonymous):

Then how would i do it?

OpenStudy (anonymous):

Well, u know what x and k are, so you can calculate y.

OpenStudy (anonymous):

x= 6 k=50 so 50(6^2) so 50*36 = 1800lbs?

OpenStudy (anonymous):

Yup, the whole idea of a constant of proportionality is that it is constant for whatever x and y.

OpenStudy (anonymous):

Oh okay. i get it. So the last question ask, which of the conic sections correctly describes this relationship for x>0? circle, hyperbola, ellipse, or half parabola?